# Efficient Attention Mechanisms for Large Language Models: A Survey

### v2

@@ 1 Introduction
Transformer-based architectures [60] have become the de-facto choice as the backbone of modern Large Language Models (LLMs). Despite their success, the standard self-attention mechanism remains a significant computational bottleneck, with quadratic time and memory complexity with respect to input sequence length. This limitation poses substantial challenges for scaling LLMs to handle increasingly long contexts with both strong performance and high efficiency.
To address this, two major directions have emerged to reduce the time and space complexity of $\mathrm{softmax}$ Attention. The first mechanism is Linear Attention [22, 52, 17, 69, 49, 71], which seeks to reduce attention complexity by reparameterizing or approximating the softmax attention as linear operations. The second candidate is Sparse Attention [8, 21, 55, 72, 33], which restricts attention computation to a subset of the full key space based on fixed or dynamic sparsity patterns. While both approaches aim to improve efficiency, they differ significantly in formulation, design choices, and hardware implications.
@@
We categorize linear attention methods into three major paradigms. First, kernelized linear attention approximates the softmax kernel with inner products in a feature space, achieving linear complexity via random feature maps [9, 41] or fixed positive mappings [22]. Second, recurrent linear attention with forgetting mechanisms introduces position-aware recurrence, enabling long-sequence modeling through data-independent [52] or data-dependent decay [17, 69], which control how past information fades over time. Third, fast-weight and meta-learning-based formulations reinterpret linear attention as a memory update process optimized online, where models such as DeltaNet [49, 71] and TTT [50, 51] incorporate fast-learning dynamics directly into the state evolution. We also examine hardware-friendly representations of linear attention—including parallel, recurrent, and chunkwise forms—highlighting their respective trade-offs in computational complexity, memory footprint, and compatibility with training or inference workflows.
@@
We classify Sparse Attention into fixed-pattern sparsity, block sparsity, and clustering-based sparsity. Fixed-pattern sparsity adopts static token-level masks such as sliding windows, dilated positions, or designated global tokens, offering simplicity and hardware-friendliness [8, 4, 64, 14]. Block sparsity selects or routes attention at block granularity, either via heuristic scoring [65, 55, 33], trainable gating [15, 72], enabling structured memory access and efficient GPU utilization. Clustering-based sparsity organizes key-value pairs using content-based or position-aware grouping methods such as k-means or LSH, facilitating semantically aware retrieval with reduced memory overhead [24, 7, 31]. Finally, we also discuss bidirectional sparse designs extend sparsity patterns to encoder-style models. These approaches differ in sparsity granularity, selection mechanism, and their alignment with hardware primitives like FlashAttention [10], and collectively represent the foundation for efficient long-context modeling in modern Transformers.
@@
There are recent efforts to integrate efficient attention mechanisms into industry-level Pretrained Language Models. These include both pure efficient architectures—such as linear attention and state-space models, and hybrid designs that combine local and global attention patterns. Models like EAGLE [38], Falcon Mamba [78] and MiniCPM4 [57] demonstrate the scalability of purely linear or sparse approaches to the multi-billion parameter scale, offering strong performance with constant-time inference. Meanwhile, hybrid models [5, 56, 32, 26, 28, 53, 66] interleave dense, sparse, and local attention to balance computational efficiency with context modeling capacity, reflecting a growing trend toward compositional, hardware-aware attention designs in modern LLMs.
@@
  • Section 5 offers an Outlook on future directions, discussing open challenges and potential advances in algorithmic and hardware-aligned research.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center , where level=1text width=15em,font=, where level=2text width=14em,font=, where level=3text width=15em,font=, where level=4text width=26em,font=, where level=5text width=20em,font=, [ Efficient Attention Mechanisms [ Linear Attention, [ Kernelized Simulation, [ Softmax Approximator, [ Linear Transformer [22], Performer [9],
RFA [41], cosFormer [42], HedgeDog [76] ] ] ] [ Forgetting Mechanism, [ Data-Independent, [ RetNet [52], Eagle [38], Lightning [45], H3 [12] ] ] [ Data-dependent, [ Mamba [17, 11], GLA [69], xLSTM [1],
GateLoop [23], HGRN [44, 46], Finch [38] ] ] ] [ In-Context Learner, [ Test-Time Regression [ FWP [49], DeltaNet [71, 70], Longhorn [29],
TTT [50], Titans [3], Atlas [2] ] ] ] [ Others, [ Element-wise Linear Attention, [ AFT [74], RWKV [37] ] ] [ Multi-Pass Linear Attention, [ ABC [40], GSA [77] ] ] [ Bidirectional Linear Attention, [ Linformer [62], Luna [34] ] ] ] ] [ Sparse Attention, [ Fixed Sparse Attention [ Local Window Attention [ Sparse Transformer [8], GPT-3 [5], StreamingLLM [64] ] ] [ Dilated Attention [ LongNet [14], LogSparse [27] ] ] ] [ Block Sparse Attention [ Block-Sparse for Prefill [ MInference [21], FlexPrefill [25],
XAttention [65], SpargeAttn [75] ] ] [ Block-Sparse for Decode [ Quest [55], DoubleSparsity [67], ReSA [54] ] ] [ Routing-based Block Sparsity [ SeerAttention [15], Landmark [36],
MoBA [33], InfLLM-v2 [57], NSA [72] ] ] ] [ Clustering Attention [ Attention by Retrieval [ RetrievalAttention [30], ClusterKV [31], MagicPIG [7] ] ] ] [ Others [ Bidirectional Sparse Attention [ BigBird [73], Routing Transformer [48],
Reformer [24], Longformer [4], ClusterFormer [61] ] ] ] ] [ Pre-trained Models, [ Efficient Models, [ Efficient Attention in LLM [ Codestral Mamba [35], Eagle [38], Finch [38],
Goose [39], Falcon Mamba [78], MiniCPM-4 [57] ] ] ] [ Hybrid Models, [ Hybridization in LLM, [ GPT-3 [5], Jamba [28], Character.AI [6],
YOCO [53], MiniMax-01 [26], Gemma-3 [56],
Command A [66], LLaMA-4 [32] ] ] ] ] ]
@@ 2.1 Kernelized Linear Attention
where $Q,K,V\in\mathbb{R}^{L\times d}$ (with $L$ the sequence length and $d$ the model dimension per head). The softmax yields weights $\propto\exp(q_{i}^{\top}k_{j})$ for query $q_{i}$ and key $k_{j}$. Kernelized Linear attention instead finds a feature mapping $\phi(\cdot)$ such that the softmax kernel is approximated by a simple dot-product in an induced feature space: $\exp(q^{\top}k)\approx\phi(q)^{\top}\phi(k)$ [59]. Given such a $\phi$, one can rewrite attention as:
@@
Linear Tranformer [22] replaces the softmax kernel with a fixed positive feature map. In practice they set $\phi(x)=\mathrm{ELU}(x)+1$. $\mathrm{ELU}(\cdot)$ is differentiable in the whole defined region, showing a better performance with naive $\mathrm{ReLU}(\cdot)$ function.
Performer [9] introduces FAVOR+ – a Random Features scheme that unbiasedly approximates the softmax kernel. It samples randomized feature maps $\phi$ so that $E[\phi(Q)\phi(K)^{\top}]=\exp(QK\top)$. This yields a provably unbiased estimator of full softmax attention using only O(N) operations. In particular, Performers use positive orthogonal random features, which reduce variance in the approximation.
Random Feature Attention [41] is a linear attention built via Random Fourier Features for the softmax kernel. Similar to Performer, RFA leverages random mapping and triangular activation to approximate $\mathrm{softmax}$. RFA further normalizes the queries and keys before random projection to reduce variance. RFA also has a variant, RFA-Gate, which adds an optional gating mechanism for recency bias.
cosFormer [42] proposes to use cosine function to approximate $\mathrm{softmax}$. Since $\cos(a+b)=\cos a\cos b-\sin a\sin b$, cosFormer decomposes the cosine re-weighted attention $S_{ij}=Q^{\prime}_{i}K^{\prime}_{j}\cos(\frac{\pi}{2}\times\frac{i-j}{M})$ into a linear attention form.
HedgeDog [76] leverages a spiky kernel $\phi(x)=\exp(Wx+b)$ since they observe that the performance gap between Transformer and Linear Transformer is due to the lack of spiky and monotonic properties. HedgeDog shows a better attention entropy and monotonicity.
@@ 2.2.1 Data-Independent Decay
Retentive Networks (RetNet) [52] introduce a retention mechanism that replaces attention with a recurrent-style update using fixed decay coefficients. In a RetNet layer, each time step $t$ maintains a state vector $s_{t}$ that aggregates past inputs with exponential forgetting. The recurrence can be written as
@@
with $\gamma\in(0,1)$ a learned decay factor (per retention head) and $k_{t}^{\top}v_{t}$ a new contribution from the current token ($v_{t}$ is a value projection of $x_{t}$, and $k_{t}$ is a key projection). The output is then obtained by a linear “query” projection: $o_{t}=q_{t},s_{t}$. Unrolling equation 3 gives an explicit formula for retention:
@@
Eagle Eagle [38] improves RWKV design with outer-product memory, which is equivalent to Linear Attention. In RWKV series, the decay factor is parameterized as $\gamma=\exp(-\exp(w))$ where $w$ is a data-independent learnable factor. In practice, both RetNet and Eagle achieve linear inference scaling with competitive performance, using fixed decays to forget old information. Empirically, RetNet uses a fixed scalar $\gamma$ per head (often each layer has multiple retention heads with different $\gamma$ values, giving a form of multi-scale decay), whereas Eagle uses learnable scalar $w$ to parameterize decay factor.
Lightning Attention [43, 45] also proposes a linear attention layer augmented with a fixed scalar decay per head to achieve length-invariant computation speed. In Lightning Attention, the hidden state is essentially $s_{t}=\lambda s_{t-1}+k_{t}^{\top}v_{t}$ for some constant $\lambda$ (with $\lambda$ learned or set by the model), which is in the same spirit as RetNet’s $\gamma$ but optimized for hardware efficiency.
H3 [12] introduces a recurrent state-space model [18] into linear attention, using learned, data-independent exponential decay via SSM for key-value outer-product hidden state. While Linear Attention enables efficient training via chunk-wise computation, H3 requires explicit state expansion for SSM computation, thus restricts the head dimension contributing to limited expressiveness.
@@ 2.2.2 Data-Dependent Decay
where $S_{t-1}$ is the previous state, and $G_{t}$ is a gating tensor determined by the token $x_{t}$ . If $G_{t}$ is close to 0 in some component, the past state in that component is largely forgotten at time $t$; if $G_{t}\approx 1$, the past is retained. Unlike the constant $\gamma$ in RetNet, here $G_{t}$ varies with $t$ via $x_{t}$. Two notable examples of this strategy in large language model design are Mamba [17, 11] and Gated Linear Attention (GLA) [69].
Mamba is a recurrent state-space model that endows the state decay rates with input dependence. In each Mamba layer, the base state evolution is similar to S4 [18], but the state matrix is effectively made dynamic. $G_{t}$ is a group-wise vector ranging from 0 to 1 as a dynamic forget gate. Which bridges a gap between attention and pure SSMs. Empirical results show that Mamba2 can outperform Transformers of similar or even larger size on language modeling tasks, highlighting the power of data-dependent decay in long-sequence modeling.
@@
Beyond these, several other models similarly endow their recurrence with content-dependent gates.
xLSTM [1] replaces the standard sigmoid forget gate with an exponential transform of linear gate signals (with normalization), yielding a smooth, input-conditioned decay on its cell state.
GateLoop [23] applies head-wise gate based on retention, which enables a simple but effective data-dependent decay while maintaining efficient hardware implementation.
HGRN [44] introduces gated recurrence in the linear RNNs. HGRN2 [46] further add state expansion into HGRN framework. State expansion is equivalent to key-value out product in Linear Attention.
Finch [38] employs data-dependent gating on Eagle. Since Eagle is similar to Retention with other orthogonal modifications, Finch also shows deep connection with above models.
@@ 2.3 Linear Attention as In-Context Learners
While large Transformer models inherently exhibit in-context learning by interpreting the prompt as a form of training data, recent innovations have integrated fast learning rules directly into the attention mechanism, effectively treating sequence processing as an online training process. FWP [49] establish a formal equivalence between existing linear attention mechanisms and Fast Weight Programmers. In the FWP paradigm, a slow neural network learns to program the "fast weights" of another network, often via additive outer products of self-invented key and value patterns. This section explores several models, including DeltaNet [71, 70], Longhorn [29], Test-Time Training (TTT) layers [50, 51], Titans [3], that exemplify this paradigm of leveraging linear attention as an in-context learner through mechanisms like fast weight updates, viewed through a meta-learning lens.
@@
Linear Attention [22] $S_{t}=S_{t-1}+k_{t}v_{t}^{\top}$
RetNet [52] $S_{t}=\gamma S_{t-1}+k_{t}v_{t}^{\top}$
GLA [69] $S_{t}=S_{t-1}\mathrm{Diag}(a_{t})+k_{t}v_{t}^{\top}$
Mamba [17] $S_{t}=\alpha_{t}S_{t-1}+b_{t}k_{t}v_{t}^{\top}$
HGRN-2 [46] $S_{t}=S_{t-1}\mathrm{Diag}(a_{t})+(1-a_{t})v_{t}^{\top}$
DeltaNet [71] $S_{t}=S_{t-1}(1-\beta_{t}k_{t}k_{t}^{\top})+\beta_{t}k_{t}v_{t}^{\top}$
Gated DeltaNet [70] $S_{t}=\alpha_{t}S_{t-1}(1-\beta_{t}k_{t}k_{t}^{\top})+\beta_{t}k_{t}v_{t}^{\top}$
TTT [51] $S_{t}=S_{t-1}-\eta_{t}\nabla_{S}\ell(S_{t-1};k_{t},v_{t})$

### v3

@@ 1 Introduction
Transformer-based architectures [transformer] have become the de-facto choice as the backbone of modern Large Language Models (LLMs). Despite their success, the standard self-attention mechanism remains a significant computational bottleneck, with quadratic time and memory complexity with respect to input sequence length. This limitation poses substantial challenges for scaling LLMs to handle increasingly long contexts with both strong performance and high efficiency.
To address this, two major directions have emerged to reduce the time and space complexity of $\mathrm{softmax}$ Attention. The first mechanism is Linear Attention [linear-transformer, retnet, mamba, gla, fastweight, deltanet], which seeks to reduce attention complexity by reparameterizing or approximating the softmax attention as linear operations. The second candidate is Sparse Attention [sparsetransformer, jiang2024minference, tang2024questqueryawaresparsityefficient, nsa, moba], which restricts attention computation to a subset of the full key space based on fixed or dynamic sparsity patterns. While both approaches aim to improve efficiency, they differ significantly in formulation, design choices, and hardware implications.
@@
We categorize linear attention methods into three major paradigms. First, kernelized linear attention approximates the softmax kernel with inner products in a feature space, achieving linear complexity via random feature maps [performer, rfa] or fixed positive mappings [linear-transformer]. Second, recurrent linear attention with forgetting mechanisms introduces position-aware recurrence, enabling long-sequence modeling through data-independent [retnet] or data-dependent decay [mamba, gla], which control how past information fades over time. Third, fast-weight and meta-learning-based formulations reinterpret linear attention as a memory update process optimized online, where models such as DeltaNet [fastweight, deltanet] and TTT [ttt, ttt2024] incorporate fast-learning dynamics directly into the state evolution. We also examine hardware-friendly representations of linear attention—including parallel, recurrent, and chunkwise forms—highlighting their respective trade-offs in computational complexity, memory footprint, and compatibility with training or inference workflows.
@@
We classify Sparse Attention into fixed-pattern sparsity, block sparsity, and clustering-based sparsity. Fixed-pattern sparsity adopts static token-level masks such as sliding windows, dilated positions, or designated global tokens, offering simplicity and hardware-friendliness [sparsetransformer, longformer, streamingllm, longnet]. Block sparsity selects or routes attention at block granularity, either via heuristic scoring [xu2025xattention, tang2024questqueryawaresparsityefficient, moba], trainable gating [seerattention, nsa], enabling structured memory access and efficient GPU utilization. Clustering-based sparsity organizes key-value pairs using content-based or position-aware grouping methods such as k-means or Local Sensitive Hashing (LSH), facilitating semantically aware retrieval with reduced memory overhead [reformer, chen2024magicpiglshsamplingefficient, liu2024clusterkvmanipulatingllmkv]. Finally, we also discuss bidirectional sparse designs extend sparsity patterns to encoder-style models. These approaches differ in sparsity granularity, selection mechanism, and their alignment with hardware primitives like FlashAttention [flashattn2], and collectively represent the foundation for efficient long-context modeling in modern Transformers.
@@
There are recent efforts to integrate efficient attention mechanisms into industry-level Pretrained Language Models. These include both pure efficient architectures—such as linear attention and state-space models, and hybrid designs that combine local and global attention patterns. Models like EAGLE [rwkv56], Falcon Mamba [falcon_mamba] and MiniCPM4 [minicpm4] demonstrate the scalability of purely linear or sparse approaches to the multi-billion parameter scale, offering strong performance with constant-time inference. Meanwhile, hybrid models [gpt3, gemma3, llama4, minimax01, jamba, yoco, rnope] interleave dense, sparse, and local attention to balance computational efficiency with context modeling capacity, reflecting a growing trend toward compositional, hardware-aware attention designs in modern LLMs.
@@
This survey is motivated by a pivotal transition in the landscape of Large Language Models. While previous development was predominantly driven by a parameter-scaling paradigm aimed at maximizing capability, the current frontier has shifted toward an efficiency-centric orientation. As demands for million-token contexts and low inference cost, the quadratic time-space complexity of canonical self-attention constitutes the principal bottleneck inhibiting scalability. Consequently, the redesign of attention mechanisms has emerged as the critical frontier for architectural innovation. Synthesizing these developments is essential to bridge the disconnect between theoretical algorithmic design and the practical latencies imposed by modern hardware accelerators, thereby laying the groundwork for the next generation of scalable and commercially viable models.
@@
  • Section 5 and 6 offers Takeaways and an Outlook on future directions, discussing open challenges and potential advances in algorithmic and hardware-aligned research.
@@ 2.1 Kernelized Linear Attention
where $Q,K,V\in\mathbb{R}^{L\times d}$ (with $L$ the sequence length and $d$ the model dimension per head). The softmax yields weights $\propto\exp(q_{i}^{\top}k_{j})$ for query $q_{i}$ and key $k_{j}$. Kernelized Linear attention instead finds a feature mapping $\phi(\cdot)$ such that the softmax kernel is approximated by a simple dot-product in an induced feature space: $\exp(q^{\top}k)\approx\phi(q)^{\top}\phi(k)$ [kernel-attn]. Given such a $\phi$, one can rewrite attention as:
@@
Linear Tranformer [linear-transformer] replaces the softmax kernel with a fixed positive feature map. In practice they set $\phi(x)=\mathrm{ELU}(x)+1$. $\mathrm{ELU}(\cdot)$ is differentiable in the whole defined region, showing a better performance than naive $\mathrm{ReLU}(\cdot)$ function.
Performer [performer] introduces FAVOR+ – a Random Features scheme that unbiasedly approximates the softmax kernel. It samples randomized feature maps $\phi$ so that $E[\phi(Q)\phi(K)^{\top}]=\exp(QK^{\top})$. This yields a provably unbiased estimator of full softmax attention using only O(N) operations. In particular, Performers use positive orthogonal random features, which reduce variance in the approximation.
Random Feature Attention [rfa] is a linear attention built via Random Fourier Features for the softmax kernel. Similar to Performer, RFA leverages random mapping and triangular activation to approximate $\mathrm{softmax}$. RFA further normalizes the queries and keys before random projection to reduce variance. RFA also has a variant, RFA-Gate, which adds an optional gating mechanism for recency bias.
cosFormer [cosformer] proposes to use cosine function to approximate $\mathrm{softmax}$. Since $\cos(a+b)=\cos a\cos b-\sin a\sin b$, cosFormer decomposes the cosine re-weighted attention $S_{ij}=Q^{\prime}_{i}K^{\prime}_{j}\cos(\frac{\pi}{2}\times\frac{i-j}{M})$ into a linear attention form. $S_{ij}$ represents the attention score between the $i$-th query and $j$-th key. $Q^{\prime}_{i}$ and $K^{\prime}_{j}$ is a non-negative projection.
HedgeDog [hedgehog] leverages a spiky kernel $\phi(x)=\exp(Wx+b)$ since they observe that the performance gap between Transformer and Linear Transformer is due to the lack of spiky and monotonic properties. HedgeDog shows a better attention entropy and monotonicity.
@@ 2.2.1 Data-Independent Decay
Retentive Networks (RetNet) [retnet] introduce a retention mechanism that replaces attention with a recurrent-style update using fixed decay coefficients. In a RetNet layer, each time step $t$ maintains a state vector $s_{t}$ that aggregates past inputs with exponential forgetting. The recurrence can be written as
@@
with $\gamma\in(0,1)$ a learned decay factor (per retention head) and $k_{t}^{\top}v_{t}$ a new contribution from the current token ($v_{t}$ is a value projection of $x_{t}$, and $k_{t}$ is a key projection). The output is then obtained by a linear “query” projection: $o_{t}=q_{t}S_{t}$. Unrolling equation 3 gives an explicit formula for retention:
@@
Eagle [rwkv56] improves RWKV design with outer-product memory, which is equivalent to Linear Attention. In RWKV series, the decay factor is parameterized as $\gamma=\exp(-\exp(w))$ where $w$ is a data-independent learnable factor. In practice, both RetNet and Eagle achieve linear inference scaling with competitive performance, using fixed decays to forget old information. Empirically, RetNet uses a fixed scalar $\gamma$ per head (often each layer has multiple retention heads with different $\gamma$ values, giving a form of multi-scale decay), whereas Eagle uses learnable scalar $w$ to parameterize decay factor.
Lightning Attention [transnormerllm, lightning] also proposes a linear attention layer augmented with a fixed scalar decay per head to achieve length-invariant computation speed. In Lightning Attention, the hidden state is essentially $s_{t}=\lambda s_{t-1}+k_{t}^{\top}v_{t}$ for some constant $\lambda$ (with $\lambda$ learned or set by the model), which is in the same spirit as RetNet’s $\gamma$ but optimized for hardware efficiency.
H3 [h3] introduces a recurrent state-space model [s4] into linear attention, using learned, data-independent exponential decay via State Space Models (SSM) for key-value outer-product hidden state. While Linear Attention supports efficient chunk-wise training, H3’s reliance on explicit state expansion for SSM computation imposes a practical bottleneck on the head dimension, thereby compromising its expressiveness.
@@ 2.2.2 Data-Dependent Decay
where $S_{t-1}$ is the previous state, and $G_{t}$ is a gating tensor determined by the token $x_{t}$ . If $G_{t}$ is close to 0 in some component, the past state in that component is largely forgotten at time $t$; if $G_{t}\approx 1$, the past is retained. Unlike the constant $\gamma$ in RetNet, here $G_{t}$ varies with $t$ via $x_{t}$. Two notable examples of this strategy in large language model design are Mamba [mamba, mamba2] and Gated Linear Attention (GLA) [gla].
Mamba is a recurrent state-space model that endows the state decay rates with input dependence. In each Mamba layer, the base state evolution is similar to S4 [s4], but the state matrix is effectively made dynamic. $G_{t}$ is a group-wise vector ranging from 0 to 1 as a dynamic forget gate. Which bridges a gap between attention and pure SSMs. Reported empirical results suggest that Mamba2 can achieve competitive or superior performance compared to Transformers of similar or larger scale on specific language modeling tasks, reflecting the potential of data-dependent decay in long-sequence modeling.
@@
Beyond these, several other models similarly endow their recurrence with content‑dependent gates.
xLSTM [xlstm] replaces the standard sigmoid forget gate with an exponential transform of linear gate signals (with normalization), yielding a smooth, input‑conditioned decay on its cell state.
GateLoop [gateloop] applies head-wise gate on retention, which enables a simple but effective data-dependent decay while maintaining efficient hardware implementation.
HGRN [hgrn] introduces gated recurrence in the linear RNNs. HGRN2 [hgrn2] further add state expansion into HGRN framework. State expansion is equivalent to key-value out product in Linear Attention.
Finch [rwkv56] employs data‑dependent gating on Eagle. Since Eagle is similar to Retention with other orthogonal modifications, Finch also shows deep connection with above models.
@@ 2.3 Linear Attention as In-Context Learners
While large Transformer models inherently exhibit in-context learning by interpreting the prompt as a form of training data, recent innovations have integrated fast learning rules directly into the attention mechanism, effectively treating sequence processing as an online training process. FWP [fastweight] establish a formal equivalence between existing linear attention mechanisms and fast weight learning. In the FWP paradigm, a slow neural network learns to program the "fast weights" of another network, often via additive outer products of self-invented key and value patterns. This section explores several models, including DeltaNet [deltanet, gdn], Longhorn [longhorn], Test-Time Training (TTT) layers [ttt, ttt2024], Titans [titans], that exemplify this paradigm of leveraging linear attention as an in-context learner through mechanisms like fast weight updates, viewed through a meta-learning lens.
@@
Linear Attention [linear-transformer] $S_{t}=S_{t-1}+k_{t}v_{t}^{\top}$
RetNet [retnet] $S_{t}=\gamma S_{t-1}+k_{t}v_{t}^{\top}$
GLA [gla] $S_{t}=S_{t-1}\mathrm{Diag}(a_{t})+k_{t}v_{t}^{\top}$
Mamba [mamba] $S_{t}=\alpha_{t}S_{t-1}+b_{t}k_{t}v_{t}^{\top}$
HGRN-2 [hgrn2] $S_{t}=S_{t-1}\mathrm{Diag}(a_{t})+(1-a_{t})v_{t}^{\top}$
DeltaNet [deltanet] $S_{t}=S_{t-1}(1-\beta_{t}k_{t}k_{t}^{\top})+\beta_{t}k_{t}v_{t}^{\top}$
Gated DeltaNet [gdn] $S_{t}=\alpha_{t}S_{t-1}(1-\beta_{t}k_{t}k_{t}^{\top})+\beta_{t}k_{t}v_{t}^{\top}$
TTT [ttt2024] $S_{t}=S_{t-1}-\beta_{t}\nabla_{S}\ell(S_{t-1};k_{t},v_{t})$